OpenStudy (anonymous):

Simplify (b^3)^4 times (b^2)^5 will give metal to best answer A. b^120 B. b^2 C. b^22 D. 2b^22

OpenStudy (anonymous):

\[(b^3)^4 \times (b^2)^5\]

OpenStudy (anonymous):

Consider: \[2^3*2^2 = 8*4 = 32 = 3^5\] and: \[(2^3)^2 = 8^2 = 64 = 3^6\] This suggests, perhaps, that: \[b^n*b^m = b^{n+m}\] and that: \[(b^n)^m = b^{nm}\] This is in fact the case. Applying these, \[(b^3)^4 \times (b^2)^5 = b^{3\times4} \times b^{2\times5} = b^{12+10} = b^{22}\] Thus the answer is C.

OpenStudy (anonymous):

ok, thanks!